SW&IT Meeting Minutes 2019-08-19
Attendees: Andrew, Carl, David, Igor, Matthieu, Petr, Stephen, Xianyu, Yang
Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.
| LR20190729 - Bec, Matthieu
| LR20190729 - Holmberg, Carl
| LR20190729 - Colson, Andrew
| LR20190729 - Edmeades, David
| LR20190729 - Sola, Igor
| LR20190729 - Kubanek, Petr
| LR20190729 - Hooper, Stephen
| LR20190729 - Zhang, Yang
| LR20190729 - Zhang, Xianyu
| LR20190812 - Bec, Matthieu
| LR20190812 - Holmberg, Carl
| LR20190812 - Colson, Andrew
| LR20190812 - Edmeades, David
| LR20190812 - Sola, Igor
| LR20190812 - Kubanek, Petr
| LR20190812 - Hooper, Stephen
| LR20190812 - Zhang, Yang
| LR20190812 - Zhang, Xianyu
News from management
Previous Action Items
- mountain AGW support
- PID tuning
- vim syntax file for PMAC code
- modifing PMAC code - automatic gather, break wait time as parameter
- OAC - cliaoc to move axes, dump gathered data
- oneliner to list used DPRAM: [pkubanek@pkubanek-dev pmac_fw]$ grep -- "->\(DP\|X\|Y\):\$6C" * | sed "s/.*\(DP\|X\|Y\):\(\$[^, \t]*\).*/\2/" | sort | uniq
- learned hard way why #pragma pack is important
- ECS patch (installed Sunday) - PLC compressor values (git cherry-pick), had to revert changes accidentally occured (wrong merge during merging to master) in 2018B
- http/GET for sharing/bookmark
- Vibration OVMS PSD, parallelized
- Misc UI cleaning
- smartsheet formula for LF
- review imonnit
- Working on LBC skyflat integration strategy (need to create a github issue).
- Change LBC camera configure and parameters (both in the conf files and lbc.zip).
- MODS scripts update on the OBS machines (in-progress discussion with OSU team).
- Working on creating new MODS telemetry data.
- Review existing setup for a meeting with Al Conrad.
- Review mongodb material for a final database scheme.
- Minor changes to some of the LBC alarm scripts (need to check if github sync is needed).
- Mountain work
- Lucix backup
- Discovered that luci.luci and lucix have different MD setups
- Have DR copy of lucix on GDrive for safekeeping
- Clone VM1 onto SSD
- Vault script finalized, worked all weekend.
- Colocation prep
- Document network info for colo team
- VM1 prep for migrating VMs
- worked on bug fixes and release
- program ID, luci filters, guide stars, MODS XSLT bug fixes
- Hot Wiring the Transient Universe VI conference
- Fixing some of the MODS scripting issues if I find the time
- Produce some version of the LUCI launch scripts so Stephen can build the 64 bit machine.
- Monnit telemetry
- Initial script rev.
- Naive approach, output looked good in hdfview, seriously sub-optimal in h5dump.
- 10 minutes between samples probably not suitable for HDF5.
- Matthieu provided JSON/Mongo-based demo, will expand on it.
- Alarm Handling System (AHS, aka CHAOS)
- Need to review data dictionary to see what systems are currently writing alert data there, after which we can determine what's missing.
- Merged tcsweb PR #11 - fix to write atomically to fire and grab radar image from http://info
- Rebuilt tcsweb rpm on f30 and x86_64
- Tested UMAC library
- Started design and flow of OAC telemetry data
- OAC telemetry
- Test FACSUM on NUC
- ARGOS software tagup?
- mountain Tuesday
- Switch stack 0202 repaired, ordered parts for 0401 and 0207
- Restarted computer rooms
- obs kickstart tweaks
- RPM for big displays.
- git on AO machines after power up
- parts search e-mail for Bonnie/Dan
- RPM for perl path
- up the mountain when parts come in
- kickstart obs VM for mountain
- room 507 upgrades
- laptop install for Tyler.